# Chiral superconductivity from parent Chern band and its non-Abelian generalization

### v1

@@ Abstract
The topological data, including ground state degeneracy, Hall conductance, anyon charge and statistics can be extracted from the standard $K$-matrix theory [84]. We find that the Hall conductance is $\sigma_{H}=0$ or $\sigma_{H}=1$ for $m$ positive or negative, corresponding to a trivial or an integer quantum Hall (IQH) phase, respectively. Meanwhile, since the absolute value of the determinant of the $K$-matrix for Eq. (10) equals to one, neither of the two phases are topological ordered due to the trivial ground state degeneracy. Thus the $\sigma_{H}=1$ ($\sigma_{H}=0$) phase is topological equivalent to the QAH crystal (gapped BEC) phase in the phase diagram Fig. [1.(a)]. Though the gapped BEC phase exhibits superconductivity while in the $\sigma_{H}=0$ phase the composite fermion does not pair, these two phases are topological equivalent as discussed in the study of proximity effect [85, 5].
When $r<0$, $\Phi$ is condensed, driving the system into superconducting phase for the composite fermion $\psi$. Correspondingly, the $U(1)$ gauge field $a$ is Higgsed to $\mathbb{Z}_{2}$. The pairing potential ensures $\psi$ to remain gapped regardless the sign of $m$, as long as $|m|<|\lambda\left<\Phi\right>|$. Thus, integrating out $\psi$ naively yields the effective Lagrangian
@@
where $\tilde{a}$ is the Lagrange multiplier introduced such that $a$ is indeed a $\mathbb{Z}_{2}$ gauge field. The Hall conductance in this phase is $\sigma_{H}=1/2$, and the vortex whose current is coupled to $\tilde{a}$ has charge $Q_{\mathrm{v}}=1/4$ and self-statistical angle $\theta_{\mathrm{v}}=\pi/8$. The quasi hole coupled to gauge field $b$ has half charge $Q_{\mathrm{h}}=1/2$ and semionic self-statistics $\theta_{\mathrm{h}}=\pi/2$. These topological data exactly match those in the non-Abelian Moore-Read quantum Hall state. However, Eq. (11) only captures the Abelian part of the anyon statistics. To incorporate non-Abelian statistics, one need to include Majorana zero modes arising naturally in the chiral superconducting phase of the composite fermion. One way to include Majorana after integrating out $\psi$ is to substitute the $\tilde{a}\mathrm{d}a$ term by $(\tilde{a}+\eta i\mathrm{d}\eta/4)\mathrm{d}a$ [86], where $\eta$ describes the Majorana zero mode trapped at the superconducting vortex core. This gives rise to the desired non-Abelian statistics [86] in the Moore-Read state. The non-Abelian Moore-Read quantum Hall phase is corresponding to $(p+ip)$-wave chiral TSC phase in Fig. [1.(a)]. It is important to note that the Pfaffian and the anti-Pfaffian quantum Hall phases with similar non-Abelian statistics can be also achieved by tuning the pairing symmetry of the composite fermion in Eq. (9[74].
In the composite fermion theory, there is an additional phase without a counterpart in the physical electron model and phase diagram Fig. [1.(a)]. To see this, note that when the composite fermion pairs, the gauge field $a$ is a $\mathbb{Z}_{2}$ gauge field which can be confined by increasing the Maxwell coupling $\kappa$ through a 3D Ising transition [87]. When $a$ is confined, the topological part of the theory reads
@@
where in passing to the second equality the gapped gauge field $b$ is integrated out, yielding a same Hall conductance $\sigma_{H}=1/2$ as the Moore-Read phase discussed previously. This phase is a CSL phase with an Abelian semion whose current is coupled to $b$. Therefore, we conjecture the existence of the CSL phase in the vicinity of the Moore-Read quantum Hall phase with Hall conductance $\sigma_{H}=1/2$ in multi-layer graphene systems.
Equipped with above analysis, we can map out the phase diagram of the composite fermion theory Eq. (9), as shown in Fig. [1 (b-d)]. Fig. [1.(b)] shows when Maxwell coupling $\kappa$ is small, there will be three phases, IQH, Moore-Read, and trivial gapped phases. This is similar to the result in physical electron model shown in Fig. [1.(a)]. When $\kappa$ is strong, the confinement of the ${\mathbb{Z}}_{2}$ gauge field will drive the Moore-Read phase into the CSL phase, as shown in Fig. [1.(d)]. Fig. [1.(c)] is plotted when the composite fermion mass $m<0$, which explictly shows the 3D Ising transition between the Moore-Read phase and the CSL phase.
Discussion– In this work, we discuss the phase diagram of spin-valley polarized electrons under parent Chern band in the presence of both attractive and repulsive interactions. While our study is focused on the parent band with Chern number equals to one, it’s natural to generalize this set up to the case with higher Chern numbers. We conjecture that more ways of asymmetric pairing, such as $(f+if)$-wave and $(h+ih)$-wave, can emerge. Our field theory analysis predicts the existence of chiral spin liquid in the vicinity of the non-Abelian Moore-Read quantum Hall phase, and it will be an interesting future direction to verify it with state-of-the-art numerical methods.
Acknowledgements– We thank Srinivas Raghu, Jing-Yuan Chen, and Zhaoyu Han for discussions on related topics. Y.-Q. W. acknowledges the support from the JQI postdoctoral fellowship at the University of Maryland. Z.-Q. G. is funded through the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC02-05-CH11231 (Theory of Materials program KC2301).
@@
I Hatree-Fock calculation for insulating phase
@@
×$\times\delta_{{\bm{k}}_{1}+{\bm{g}}_{1}+{\bm{k}}_{2}+{\bm{g}}_{2}-{\bm{k}}_{3}-{\bm{g}}_{3}-{\bm{k}}_{4}-{\bm{g}}_{4}}.$
Under the Hartree-Fock approximation, the interaction Hamiltonian can be decomposed into the Hartree term with ${\bm{k}}_{1}={\bm{k}}_{4}$ and ${\bm{k}}_{2}={\bm{k}}_{3}$:
$H_{\mathrm{H}}=E_{\mathrm{H}}+\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}_{1},{\bm{g}}_{2},{\bm{g}}_{3},{\bm{g}}_{4}}$ $2S({\bm{g}}_{1}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{1}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{2}+{\bm{g}}_{3}}P_{{\bm{g}}_{1},{\bm{g}}_{4}}({\bm{k}}_{1})c^{\dagger}_{{\bm{k}}_{2}+{\bm{g}}_{2}}c_{{\bm{k}}_{2}+{\bm{g}}_{3}}$ (S3)
×$\times\delta_{{\bm{g}}_{1}+{\bm{g}}_{2}-{\bm{g}}_{3}-{\bm{g}}_{4}},$
@@
$H_{\mathrm{F}}=E_{\mathrm{F}}-\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}_{1},{\bm{g}}_{2},{\bm{g}}_{3},{\bm{g}}_{4}}$ $2S({\bm{k}}_{1}+{\bm{g}}_{1}-{\bm{k}}_{2}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{2}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{1}+{\bm{g}}_{3}}P_{{\bm{g}}_{1}{\bm{g}}_{3}}({\bm{k}}_{1})c^{\dagger}_{{\bm{k}}_{2}+{\bm{g}}_{2}}c_{{\bm{k}}_{2}+{\bm{g}}_{4}}$ (S4)
×$\times\delta_{{\bm{g}}_{1}+{\bm{g}}_{2}-{\bm{g}}_{3}-{\bm{g}}_{4}}$
@@
In calculation we choose total 7 BZs with the first BZ centered and the other 6 BZs packed around it, and 108 ${\bm{k}}$ points within each BZ. The electron filling is set to be $\nu=1$, which means the total number of electron $N$ is equal to the number of ${\bm{k}}$ points in the first BZ, i.e. $N=108$. The ground state energy of the insulating phase under the Hartree-Fock approximation is
$E_{\mathrm{HF}}=\sum_{{\bm{k}}\in\mathrm{BZ}}\bra{{\bm{k}}}H_{0}+H_{\mathrm{H}}+H_{\mathrm{F}}\ket{{\bm{k}}}=-E_{\mathrm{H}}-E_{\mathrm{F}}+\sum_{{\bm{k}}\in\mathrm{BZ}}\epsilon({\bm{k}})$ (S8)
@@ II Mean-field calculation for superconducting phase
which respects spatial asymmetry $\Delta(-{\bm{k}})=-\Delta({\bm{k}})$ for the pairing of spinless fermion. Then the BdG Hamiltonian reads
@@
III $K$-matrix calculation and non-Abelian statistics in field theory
As stated in the main text, in the non-superconducting phases of the composite fermion, the topological part of the Lagragian reads
@@ III $K$-matrix calculation and non-Abelian statistics in field theory
which does not host topological order since $|\det K|=1$. The charge vector is $\mathbf{p}=(0,1)^{\mathbf{T}}$, and hence the Hall conductance is $\sigma_{H}=\mathbf{p}^{\mathbf{T}}K_{\mathrm{QH}}^{-1}\mathbf{p}=-\Theta(-m)$. Thus $\sigma_{H}=0$ for $m>0$ and $\sigma_{H}=1$ for $m<0$. A topological phase transition occur at $m=0$, where the composite fermiom is gapless.
In the Moore-Read phase, the effective Lagrangian capturing the Abelian part of the anyon statistics is
@@
with charge vector $\mathbf{q}=(0,0,1)^{\mathbf{T}}$. The Hall conductance in this phase $\sigma_{H}=\mathbf{q}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{q}=1/2$. The vortex carrying “magnetic charge” under the $\mathbb{Z}_{2}$ gauge field $a$ has its current coupled to $\tilde{a}$. Thus, its $l$-vector is $\mathbf{l}_{\mathrm{v}}=(0,1,0)$. The vortex carries charge $Q_{\mathrm{v}}=\mathbf{l}_{\mathrm{v}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{q}=1/4$ and self-statistical angle $\theta_{\mathrm{v}}=\pi\mathbf{l}_{\mathrm{v}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{l}_{\mathrm{v}}=\pi/8$. The quasi hole whose current is coupled to gauge field $c$ has $\mathbf{l}_{\mathrm{h}}=(0,0,1)$. Therefore it has half charge $Q_{\mathrm{h}}=\mathbf{l}_{\mathrm{h}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{q}=1/2$ and semionic self-statistics $\theta_{\mathrm{h}}=\pi\mathbf{l}_{\mathrm{h}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{l}_{\mathrm{h}}=\pi/2$. The incorporation of Majorana zero modes reads
@@
where $j^{\mathrm{v}}_{\mu}$ and $j^{\mathrm{h}}_{\mu}$ are the currents of the vortices and the quasi holes, respectively, while the Majorana zero modes are described by $\eta$. Equations of motion for $\tilde{a}$ and $b$ gives rise to
@@
after integrating out all the dynamical gauge fields. Here $1/\partial$ is a shorthand of the current-current interaction mediated by gauge fields. The first term describing vortices trapping Majorana zero modes gives rise to the desired non-Abelian statistics. The charge and the Abelian self-statistical angle of vortices, $Q_{\mathrm{v}}=1/4$ and $\theta_{\mathrm{v}}=\pi/8$ respectively, can also be manifestly read off, and similarly for the quasi holes.

### v3

@@ Abstract
The topological data, including ground state degeneracy, Hall conductance, anyon charge and statistics can be extracted from the standard $K$-matrix theory [89]. We find that the Hall conductance is $\sigma_{H}=0$ or $\sigma_{H}=1$ for $m$ positive or negative, corresponding to a trivial or an integer quantum Hall (IQH) phase, respectively. Meanwhile, since the absolute value of the determinant of the $K$-matrix for Eq. (9) equals to one, neither of the two phases are topological ordered due to the trivial ground state degeneracy. Thus the $\sigma_{H}=1$ ($\sigma_{H}=0$) phase is topological equivalent to the AHC (gapped BEC) phase in the phase diagram Fig. [1.(a)]. Though the gapped BEC phase exhibits superconductivity while in the $\sigma_{H}=0$ phase the composite fermion does not pair, these two phases are topological equivalent as discussed in the study of proximity effect [90, 5].
[…]
@@
where $\tilde{a}$ is the Lagrange multiplier introduced such that $a$ is indeed a $\mathbb{Z}_{2}$ gauge field  [^4] Integrating out Majorana fermion will generate a gravitational CS term [76], which is not the focus of our study here and omitted in the field theory Eq. (10). Self CS term of the $\mathbb{Z}_{2}$ gauge field is not generated since Majorana fermion does not suffer the parity anomaly.. The Hall conductance in this phase is $\sigma_{H}=1/2$, and the vortex whose current is coupled to $\tilde{a}$ has charge $Q_{\mathrm{v}}=1/4$ and self-statistical angle $\theta_{\mathrm{v}}=\pi/8$. The quasi hole coupled to gauge field $b$ has half charge $Q_{\mathrm{h}}=1/2$ and semionic self-statistics $\theta_{\mathrm{h}}=\pi/2$. These topological data exactly match those in the nonabelian Moore–Read quantum Hall state. However, Eq. (10) only captures the Abelian part of the anyon statistics. To incorporate nonabelian statistics, one needs to include Majorana zero modes arising naturally in the chiral superconductor phase of the composite fermion. One way to include Majorana after integrating out $\psi$ is to substitute the $\tilde{a}\mathrm{d}a$ term by $(\tilde{a}+\eta i\mathrm{d}\eta/4)\mathrm{d}a$ [92], where $\eta$ describes the Majorana zero mode trapped at the vortex core. This gives rise to the desired nonabelian statistics [92] in the Moore–Read state. The nonabelian Moore–Read quantum Hall phase is corresponding to $(p+ip)$-wave chiral TSC phase in Fig. [1.(a)]. It is important to note that the Pfaffian and the anti-Pfaffian quantum Hall phases with similar nonabelian statistics can be also achieved by tuning the pairing symmetry of the composite fermion in Eq. (8[76].
In the composite fermion theory, there is an additional phase without a counterpart in the physical electron model and phase diagram Fig. [1.(a)]. To see this, note that when the composite fermion pairs, the gauge field $a$ is a $\mathbb{Z}_{2}$ gauge field which can be confined by increasing the Maxwell coupling $\kappa$ through a confinement-deconfinement transition [93]. When $a$ is confined, the topological part of the theory reads
@@
where in passing to the second equality the gapped gauge field $b$ is integrated out, yielding a same Hall conductance $\sigma_{H}=1/2$ as the Moore–Read phase discussed previously. This phase is a CSL phase with an Abelian semion whose current is coupled to $b$. Therefore, we conjecture the existence of the CSL phase in the vicinity of the Moore–Read quantum Hall phase with Hall conductance $\sigma_{H}=1/2$ in multilayer graphene systems.
Equipped with above analysis, we can map out the phase diagram of the composite fermion theory Eq. (8), as shown in Fig. [1 (b-d)]. Fig. [1.(b)] shows when Maxwell coupling $\kappa$ is small, there will be three phases, IQH, Moore–Read, and trivial gapped phases. This is similar to the result in physical electron model shown in Fig. [1.(a)]. When $\kappa$ is strong, the confinement of the ${\mathbb{Z}}_{2}$ gauge field will drive the Moore–Read phase into the CSL phase, as shown in Fig. [1.(d)]. In (b) and (d) the phase boundary between the pairing (Moore–Read or CSL) and nonpairing (IQH or trivial) phase of composite fermion should be located at $m=\pm\lambda\left<\Phi\right$> for $r<0$. Fig. [1.(c)] is plotted when the composite fermion mass $m<0$, which explicitly shows the confinement-deconfinement transition between the CSL and the Moore–Read phases. Their phase boundary can in principle depend on $\kappa$.
Discussion.– Topological BCS–BEC transitions have recently gained attention in TMD materials [85]. In tetralayer graphene, chiral superconductivity has been reported near the finite-temperature BCS–BEC crossover, although the system still predominantly resides on the BCS side [28]. This is consistent with our zero-temperature results for the BCS–BEC transition. We therefore propose that such a topological BCS–BEC transition could be detected experimentally in the future, for example via ARPES measurements that track the closing and reopening of the spectral gap [94, 95, 96].
Our field theory analysis predicts that the nonabelian Moore–Read quantum Hall phase is neighbored to a chiral pseudospin liquid phase across a confinement-deconfinement transition of $\mathbb{Z}_{2}$ gauge field. Thus, our results suggest that the chiral (pseudo-)spin liquid state may arise in the vicinity of the Moore–Read state in experiments, and vice versa. The field theory provides a solid phase diagram, closely tied to our Hartree–Fock analysis. These shed light on the search for exotic topological phases as nonabelian quantum Hall states and quantum spin liquids in future numerical simulations and experimental realization.
Acknowledgements— We acknowledge Srinivas Raghu, Jing-Yuan Chen, and Zhaoyu Han for discussions on related topics. We also thank Joel E. Moore, Zhengguang Lu, Oskar Vafek, Chunxiao Liu, Zhihuan Dong, Taige Wang, Yang-Zhi Chou and Jihang Zhu for helpful discussions. Y.-Q. W. is supported by JQI postdoctoral fellowship at the University of Maryland. Z.-Q. G. is funded through the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC02-05-CH11231 (Theory of Materials program KC2301).
[…]
@@
I Hartree–Fock calculation for insulating phase
@@
$\times\delta_{{\bm{k}}_{1}+{\bm{g}}_{1}+{\bm{k}}_{2}+{\bm{g}}_{2}-{\bm{k}}_{3}-{\bm{g}}_{3}-{\bm{k}}_{4}-{\bm{g}}_{4}}.$
Under the Hartree–Fock approximation, the interaction Hamiltonian can be decomposed into the Hartree term with ${\bm{k}}_{1}={\bm{k}}_{4}$ and ${\bm{k}}_{2}={\bm{k}}_{3}$:
$H_{\mathrm{H}}=E_{\mathrm{H}}+\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}_{1},{\bm{g}}_{2},{\bm{g}}_{3},{\bm{g}}_{4}}$ $2S({\bm{g}}_{1}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{1}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{2}+{\bm{g}}_{3}}P_{{\bm{g}}_{1},{\bm{g}}_{4}}({\bm{k}}_{1})c^{\dagger}_{{\bm{k}}_{2}+{\bm{g}}_{2}}c_{{\bm{k}}_{2}+{\bm{g}}_{3}}\delta_{{\bm{g}}_{1}+{\bm{g}}_{2}-{\bm{g}}_{3}-{\bm{g}}_{4}},$ (S3)
@@
$H_{\mathrm{F}}=E_{\mathrm{F}}-\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}_{1},{\bm{g}}_{2},{\bm{g}}_{3},{\bm{g}}_{4}}$ $2S({\bm{k}}_{1}+{\bm{g}}_{1}-{\bm{k}}_{2}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{2}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{1}+{\bm{g}}_{3}}P_{{\bm{g}}_{1}{\bm{g}}_{3}}({\bm{k}}_{1})c^{\dagger}_{{\bm{k}}_{2}+{\bm{g}}_{2}}c_{{\bm{k}}_{2}+{\bm{g}}_{4}}\delta_{{\bm{g}}_{1}+{\bm{g}}_{2}-{\bm{g}}_{3}-{\bm{g}}_{4}}.\quad$ (S4)
@@
In calculation we choose a total of 7 BZs with the first BZ centered and the other 6 BZs packed around it, and 108 $\bm{k}$} points within each BZ. The electron filling is set to be $\nu=1$, which means the total number of electron NN is equal to the number of $\bm{k}$} points in the first BZ, i.e. $N=108$. The ground state energy of the insulating phase under the Hartree–Fock approximation is
[…]
@@ II Mean-field calculation for superconducting phase
which respects spatial asymmetry $\Delta(-{\bm{k}})=-\Delta({\bm{k}})$ for the pairing of spinless fermions. Then the BdG Hamiltonian reads
@@
In numerics we check the convergence of particle-particle and particle-hole channels for each pair of $(u,v)$, and if both channels converge the energetically favored one is picked.
Note that the electron-phonon coupling potential we adopt, $U({\bm{q}})=q_{0}^{2}/(|{\bm{q}}|^{2}+q_{0}^{2})$, interpolates between the BCS limit and the case for real materials as $q_{0}$ varies. Here we would like to show that the existence of TSC is qualitatively insensitive to the details of the electron-phonon coupling potential. For $q_{0}\gg|{\bm{K}}|$, $U({\bm{q}})$ becomes nearly uniform, which is close to the BCS limit. On the contrary, when $q_{0}\ll|{\bm{K}}|$, $U({\bm{q}})$ asymptotes a delta function, which better describes materials. To see the variation of $q_{0}$ does not eliminate the chiral topological superconductor phase, we calculate the phase diagrams with $0\leq u\leq 5\epsilon_{K}$ and $0\leq v\leq 10\epsilon_{K}$ as shown in Fig. [S1 (a-c)] with (a) $q_{0}=0.05|{\bm{K}}|$, (b) $q_{0}=0.3|{\bm{K}}|$ (the same as that in the main text), and (c) $q_{0}=20|{\bm{K}}|$. It can be seen that the chiral topological superconducting phase indeed does not vanish, though disfavored by large $q_{0}$ where the uniformity of $U({\bm{q}})$ effectively leads to an enhanced coupling strength of the attractive interaction and leads to a trivial gapped BEC phase.
Figure S1: The phase diagrams with $0\leq u\leq 5\epsilon_{K}$ and $0\leq v\leq 10\epsilon_{K}$ for (a) $q_{0}=0.05|{\bm{K}}|$, (b) $q_{0}=0.3|{\bm{K}}|$, and (c) $q_{0}=20|{\bm{K}}|$. The insets zoom in the left bottom corners of the phase diagrams with higher numerical resolution.
III $K$-matrix calculation and nonabelian statistics in field theory
As stated in the main text, in the nonpairing phases of the composite fermion, the topological part of the Lagragian reads
@@ III $K$-matrix calculation and nonabelian statistics in field theory
which does not host topological order since $|\det K|=1$. The charge vector is $\mathbf{p}=(0,1)^{\mathbf{T}}$, and hence the Hall conductance is $\sigma_{H}=\mathbf{p}^{\mathbf{T}}K_{\mathrm{QH}}^{-1}\mathbf{p}=-\Theta(-m)$. Thus $\sigma_{H}=0$ for $m>0$ and $\sigma_{H}=1$ for $m<0$. A topological phase transition occurs at $m=0$, where the composite fermion is gapless.
In the Moore–Read phase, the effective Lagrangian capturing the Abelian part of the anyon statistics is
@@
with charge vector $\mathbf{q}=(0,0,1)^{\mathbf{T}}$. The Hall conductance in this phase $\sigma_{H}=\mathbf{q}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{q}=1/2$. The vortex carrying “magnetic charge” under the $\mathbb{Z}_{2}$ gauge field $a$ has its current coupled to $\tilde{a}$. Thus, its $l$-vector is $\mathbf{l}_{\mathrm{v}}=(0,1,0)$. The vortex carries charge $Q_{\mathrm{v}}=\mathbf{l}_{\mathrm{v}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{q}=1/4$ and self-statistical angle $\theta_{\mathrm{v}}=\pi\mathbf{l}_{\mathrm{v}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{l}_{\mathrm{v}}=\pi/8$. The quasi hole whose current is coupled to gauge field $b$ has $\mathbf{l}_{\mathrm{h}}=(0,0,1)$. Therefore it has half charge $Q_{\mathrm{h}}=\mathbf{l}_{\mathrm{h}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{q}=1/2$ and semionic self-statistics $\theta_{\mathrm{h}}=\pi\mathbf{l}_{\mathrm{h}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{l}_{\mathrm{h}}=\pi/2$. The incorporation of Majorana zero modes reads
@@
where $j^{\mathrm{v}}_{\mu}$ and $j^{\mathrm{h}}_{\mu}$ are the currents of the vortices and the quasi holes, respectively, while the Majorana zero modes are described by $\eta$. Equations of motion for $\tilde{a}$ and $b$ give rise to
@@
after integrating out all the dynamical gauge fields. Here $1/\partial$ is a shorthand of the current-current interaction mediated by gauge fields. The first term describing vortices trapping Majorana zero modes gives rise to the desired nonabelian statistics. The charge and the Abelian self-statistical angle of vortices, $Q_{\mathrm{v}}=1/4$ and $\theta_{\mathrm{v}}=\pi/8$ respectively, can also be manifestly read off, and similarly for the quasi holes.